OpenStudy (anonymous):

Area of a triangle area is simple to find but I am confused on this one sense I got it wrong the shape is a triangle one side of the triangle is b=4 this side is one of the longer ones there are two same length and then the bottom one , and then it says h=7 the h is down the middle of the triangle with dotted lines the area is not 28 so how do I find my answer what is the formula?

OpenStudy (anonymous):

A=1/2*b*h=1/2*4*7=14

OpenStudy (anonymous):

okay does the 1/2 change depending on the shape or does this formula adapt to any problem of this type?

OpenStudy (anonymous):

if you have the height and the base, its always the same. sometimes you have to use the pythagorean theorem or some trig to find the height

OpenStudy (anonymous):

hmm ok thank you